What is the square root of 16x divided by 9x
Answer by ptaylor(2198)

I'm not sure I understand the problem. Is it?:
(a) sqrt((16x)/9x)) or:
(b) (sqrt(16x))/(9x)
We'll do both :
(a) sqrt ((16x)/(9x))=(sqrt(16/9)(x/x))=sqrt((16/9)(1))
sqrt(16)/sqrt(9)=4/3
(b) (sqrt(16x))/(9x)=(sqrt(16))(sqrt(x))/9x=4sqrt(x)/9x=
4x^(1/2)/9x^(2/2)=4/9x^(1/2)
Hope this helps------Happy holidays.-----ptaylor